Haswell Enterprises' bonds have a 10-year maturity, a 6.25% semiannual coupon, and a par value of $1,000. The going interest rate (rd) is 5.75%, based on semiannual compounding. What is the bond's price?
Answer #1

Face Value = 1000 Coupon = 6-25°X1000 = 31.25 ylim = 5.75%. = 2.8*5*1 Time = 10x2= 20 periods. 2 Using financial calculator,